OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11

OpenStudy (anonymous):

multiply the first equation by a number which will make the term in x = 6x

OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the y's to add out? 3x - y = 3 -2x + 2y = 6 so this would be 2?

OpenStudy (anonymous):

right on

OpenStudy (anonymous):

woot! thanx

OpenStudy (anonymous):

so for the first set of equations you multiply by 3 - dont forget to multiply every term in the equation